# Pychop: Emulating Low-Precision Arithmetic in Numerical Methods and Neural Networks

### v2

@@ 3.3. Fixed-point arithmetic
Fixed-point representation provides a simpler alternative to floating-point by fixing the position of the binary point within a binary number. This conversion from floating point numbers to fixed-point numbers employs a fixed scaling factor (implicit or explicit), enabling fractional values to be represented as integers scaled by a constant, such as $2^{-f}$, where $f$ is the number of fractional bits.
@@ 3.3.2. Relation to and Deep Learning Deployment
The adoption of fixed-point quantization in deep learning offers several benefits and significant contributions that enhance its utility across various applications, and plays a pivotal role in neural networks by optimizing both inference and training phases. By leveraging fixed-point arithmetic, which uses fewer bits—such as 16-bit $Q8.8$ representations compared to 32-bit floats—this technique substantially reduces the memory footprint and accelerates multiply-accumulate (MAC) operations, which are fundamental to DNN computation. Its compatibility with hardware is a key advantage, as many edge devices like microcontrollers natively support fixed-point operations, eliminating the need for floating-point emulation and thereby boosting performance. With careful selection of the integer ($m$) and fractional ($f$) bit allocations, fixed-point quantization maintains accuracy close to that of floating-point models, a capability validated by research (Lin et al., 2016b). Ultimately, this technique has enabled the deployment of sophisticated DNNs on resource-limited platforms, significantly broadening the practical impact of AI in fields like mobile computing, real-time image processing, and autonomous navigation.
In the context of neural networks and deep learning, fixed-point quantization is applied to weights, activations, and sometimes gradients to optimize computation and storage, making it a cornerstone for efficient deployment and training. For inference, it replaces costly floating-point operations with fixed-point arithmetic, which is natively supported by many embedded systems. In training, quantization-aware training ensures the model adapts to reduced precision, minimizing accuracy loss. Fully fixed-point training, though less common, has been studied for end-to-end optimization on fixed-point hardware. This compatibility with hardware acceleration—particularly DSPs and FPGAs, which often lack native floating-point units or optimize for fixed-point operations—bridges the gap between complex DNNs (e.g., convolutional neural networks or Transformers) and practical, low-power deployment. Fixed-point quantization thus reduces memory and computational costs while enabling practical deployment of DNNs on edge devices, as explored in foundational works like (Lin et al., 2016a).
@@ 3.4. Integer arithmetic
    5X_inv = ch.dequantize(X_q) # back to floating point numbers
@@ 3.6. Seamless PyTorch / JAX Integration
In the example below, we generate three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array. By specifying the appropriate backend, we can configure Pychop to use 5 exponent bits and 10 significand bits with round to nearest mode to process these inputs accordingly. For instance, selecting the “torch” backend allows Pychop to handle X_th, the PyTorch tensor. Additionally, GPU deployment can be enabled by transferring the PyTorch tensor or JAX array to a GPU device, such as with X_th.to(’cuda’).
@@ 3.7.1. Principle and basic usage
Specify floating point precision.
@@
1optimizer = QuantizedSGD(model.parameters(), lr=0.01, momentum=0.9, exp_bits=5, sig_bits=10, rmode=1),
@@
3optimizer = QuantizedRMSProp(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=5))
@@
5optimizer = QuantizedAdagrad(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=4)
@@
7optimizer = QuantizedAdam(model.parameters(), lr=0.001, exp_bits=5, sig_bits=10, rmode=6)
Customized Low precision optimization.
@@
3quantizer = LightChop(exp_bits=5, sig_bits=10, rmode=1) # define your customized fixed-point or floating point format
@@
Post quantization.
@@ 4.1. Environmental settings
We run experiments on a Dell PowerEdge R750xa server [^5] https://front.convergence.lip6.fr/ with 2 TB of memory, Intel Xeon Gold 6330 processors (56 cores, 112 threads, 2.00 GHz), and an NVIDIA A100 GPU (80 GB HBM2, PCIe), providing robust computational power for large-scale simulations and deep learning tasks. We simulate the code in Python 3.10 and MATLAB R2024b. All simulations are performed on a single CPU and GPU.
Our analysis of image classification and object detection will primarily revolve around the following datasets:
  • MNIST(Deng, 2012): This dataset consists of handwritten digits ranging from 0 to 9, represented as grayscale images of size $28\times 28$ pixels. The dataset is widely regarded as a benchmark for image classification tasks and optical character recognition models. Each image is centered and normalized within a fixed-size frame, ensuring consistency across samples. The dataset contains 10 classes, corresponding to the 10 numerical digits, and exhibits variations in handwriting styles, stroke thickness, and digit positioning.
  • Fashion-MNIST(Xiao et al., 2017): The dataset consists of grayscale images, each of size $28\times 28$ pixels, depicting various fashion items such as clothing, footwear, and accessories. The dataset comprises 10 distinct classes, with each class representing a specific category of fashion products. The images are derived from the Zalando dataset and are designed as a more complex alternative to the MNIST dataset, introducing greater variability in shape and texture while preserving the same structural properties for benchmarking classification algorithms.
@@
  • Caltech101(Li et al., 2022): The dataset contains $224\times 224$ RGB images of objects spanning 101 diverse categories, including animals, vehicles, and household items, with an additional background class. Collected by the California Institute of Technology, it presents a challenging benchmark for image classification due to its high intraclass variability and imbalanced sample sizes, ranging from 40 to 800 images per category, offering a robust testbed for evaluating generalization across heterogeneous visual patterns.
@@
  • OxfordIIITPet(Parkhi et al., 2012): The dataset contains $224\times 224$ RGB images of pet animals across 37 breeds of cats and dogs, sourced from a collaboration between the University of Oxford and IIT. Designed to assess fine-grained classification, it features approximately 200 images per breed, split into training/validation and test sets, with significant variability in pose, lighting, and background, making it an ideal resource for studying detailed visual discrimination in real-world scenarios.
  • COCO(Lin et al., 2014): The COCO dataset comprises high-resolution color images of everyday scenes featuring objects from 80 distinct categories, including people, animals, vehicles, and household items. Originating from the Microsoft Common Objects in Context initiative, it offers a robust challenge beyond simpler datasets like ImageNet, with complex backgrounds, multiple objects per image, and annotations for both bounding boxes and segmentation masks. In our simulation code, we specifically utilize the COCO val2017 subset, which includes approximately 5,000 images from the validation split. This choice enables efficient evaluation of our quantized Faster R-CNN model’s performance on a diverse, well-annotated set without the computational overhead of the full training set (over 118,000 images) or the restricted access of the test set, ensuring rapid experimentation and reliable benchmarking of detection accuracy and latency trade-offs.
@@ 4.3.1. Image classifications
Figure 3. Image classification on Caltech101 (Custom(5, 5)).
Figure 4. Image classification on Caltech101 (bfloat16).
Figure 5. Image classification on Caltech101 (tf32).
Figure 6. Image classification on Caltech101 (fp32).
@@ 4.3.2. Object detection
Table 9. mAP@0.5:0.95 by Floating Point Representation and Rounding Method
@@
Figure 7. Object detection using bfloat16 (Red: Ground Truth, Green: Predictions).
Figure 8. Object detection tf32 (Red: Ground Truth, Green: Predictions).
Figure 9. Object detection using fp32 (Red: Ground Truth, Green: Predictions).
@@ 5. Conclusion and future work
In this work, we present Pychop, an open-source software designed as an efficient precision emulation tool for numerical methods and deep learning research. By integrating seamlessly with automatic differentiation frameworks such as PyTorch, JAX, and NumPy, Pychop bridges theoretical exploration and practical deployment, enhancing accessibility and usability in computational science. Its flexibility, comprehensive rounding support, and rigorous validation establish it as a vital resource for advancing mixed-precision numerical algorithms and deep learning applications.
To evaluate its adaptability and real-world utility, we employed Pychop to simulate post-quantization effects in image classification and object detection tasks across established datasets. These experiments provide insights into the optimal bitwidths for exponents and significands required for high-quality inference, revealing performance trade-offs to inform the selection of efficient floating-point representations for specific use cases.
Pychop is poised for continuous development and innovation. While currently focused on neural network quantization and mixed-precision computing, we plan to extend compatibility to additional frameworks, such as TensorFlow, to amplify its impact and adoption. We invite contributions from the research and development community—via GitHub pull requests and issue submissions—to enhance its functionality, optimize performance, and explore novel applications, ensuring Pychop remains a robust and versatile tool for future advancements.

### v4

@@ 3.3. Fixed-point arithmetic
Fixed-point representation provides a simpler alternative to floating-point by fixing the position of the binary point within a binary number. This conversion from floating-point numbers to fixed-point numbers employs a fixed scaling factor (implicit or explicit), enabling fractional values to be represented as integers scaled by a constant, such as $2^{-f}$, where $f$ is the number of fractional bits.
@@ 3.3.2. Relation to and Deep Learning Deployment
The adoption of fixed-point quantization in deep learning offers several benefits and significant contributions that enhance its utility across various applications and plays a pivotal role in neural networks by optimizing both inference and training phases. By leveraging fixed-point arithmetic, which uses fewer bits—such as 16-bit $Q8.8$ representations compared to 32-bit floats—this technique substantially reduces the memory footprint and accelerates multiply-accumulate (MAC) operations, which are fundamental to DNN computation. Its compatibility with hardware is a key advantage, as many edge devices like microcontrollers natively support fixed-point operations, eliminating the need for floating-point emulation and thereby boosting performance. With careful selection of the integer ($m$) and fractional ($f$) bit allocations, fixed-point quantization maintains accuracy close to that of floating-point models, a capability validated by research (Lin et al., 2016b). Ultimately, this technique has enabled the deployment of sophisticated DNNs on resource-limited platforms, significantly broadening the practical impact of AI in fields like mobile computing, real-time image processing, and autonomous navigation.
In the context of neural networks and deep learning, fixed-point quantization is applied to weights, activations, and sometimes gradients to optimize computation and storage, making it a cornerstone for efficient deployment and training. For inference, it replaces costly floating-point operations with fixed-point arithmetic, which is natively supported by many embedded systems. In training, quantization-aware training ensures the model adapts to reduced precision, minimizing accuracy loss. Fully fixed-point training, though less common, has been studied for end-to-end optimization on fixed-point hardware. This compatibility with hardware acceleration—particularly DSPs and FPGAs, which often lack native floating-point units or optimize for fixed-point operations—bridges the gap between complex DNNs (e.g., convolutional neural networks or Transformers) and practical, low-power deployment. Fixed-point quantization thus reduces memory and computational costs while enabling the practical deployment of DNNs on edge devices, as explored in foundational works like (Lin et al., 2016a).
@@ 3.4. Integer arithmetic
    5X_inv = ch.dequantize(X_q) # back to floating-point numbers
@@ 3.6. Seamless PyTorch / JAX Integration
In the example below, there are three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array. By specifying the appropriate backend, we can configure Pychop to use 5 exponent bits and 10 significand bits with round to nearest mode to process these inputs accordingly. For instance, selecting the “torch” backend allows Pychop to handle X_th, the PyTorch tensor. Additionally, GPU deployment can be enabled by transferring the PyTorch tensor or JAX array to a GPU device, such as with X_th.to(’cuda’).
@@ 3.7.1. Principle and basic usage
Specify floating-point precision.
@@
1optimizer = QuantizedSGD(model.parameters(), lr=0.01, momentum=0.9, exp_bits=5, sig_bits=10, rmode=1, **kwargs),
@@
3optimizer = QuantizedRMSProp(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=5, **kwargs)
@@
5optimizer = QuantizedAdagrad(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=4, **kwargs)
@@
7optimizer = QuantizedAdam(model.parameters(), lr=0.001, exp_bits=5, sig_bits=10, rmode=6, **kwargs)
Customized Low precision optimization for neural network deployment.
@@
3quantizer = LightChop(exp_bits=5, sig_bits=10, rmode=1) # define your customized fixed-point or floating-point format
@@
Post quantization for neural network deployment.
@@ 4.1. Environmental settings
Our experiments are simulated on a Dell PowerEdge R750xa server [^5] https://front.convergence.lip6.fr/ with 2 TB of memory, Intel Xeon Gold 6330 processors (56 cores, 112 threads, 2.00 GHz), and an NVIDIA A100 GPU (80 GB HBM2, PCIe), providing robust computational power for large-scale simulations and deep learning tasks. We simulate the code in Python 3.10 and MATLAB R2024b. All simulations are performed on a single CPU and GPU.
We simulated low-precision quantization using Pychop on a variety of benchmark datasets intended for a broad range of computer vision tasks to evaluate the effect of low-precision quantization on object recognition and image classification:
  • MNIST (Deng, 2012): The dataset comprises 60,000 training and 10,000 test grayscale images of handwritten digits (0–9), each with a resolution of $28\times 28$ pixels. Widely used as a benchmark for image classification and optical character recognition, the images are preprocessed to be centered and normalized for consistent sizing and intensity.
  • Fashion-MNIST (Xiao et al., 2017): The dataset contains 60,000 training and 10,000 test grayscale images of fashion items from Zalando’s inventory, each with a resolution of $28\times 28$ pixels. Spanning 10 classes (e.g., clothing, shoes, bags), it serves as a more complex alternative to MNIST for benchmarking image classification models.
@@
  • Caltech101 (Li et al., 2022): The dataset comprises approximately 9,144 RGB images of objects across 101 categories (e.g., animals, vehicles, household items) and an additional background class. Image sizes vary, typically around 300 pixels on the longer side, and are often resized (e.g., to $224\times 224$) for specific tasks. With imbalanced sample sizes (40–800 images per category), it provides a challenging benchmark for image classification due to its diverse and heterogeneous visual patterns.
@@
  • Oxford-IIIT Pet (Parkhi et al., 2012): The dataset contains approximately 7,349 RGB images of cats and dogs across 37 breeds (12 cats, 25 dogs), with about 200 images per breed, split into training/validation (3,680 images) and test (3,669 images) sets. This dataset supports fine-grained classification, with significant variability in pose, lighting, and background, making it suitable for real-world visual discrimination tasks. Here, images are resized to $256\times 256$ followed by a crop to $224\times 224$ for analysis.
  • COCO (Lin et al., 2014): The dataset contains RGB images across 80 categories (e.g., people, animals, vehicles, household items), designed for object detection, segmentation, and captioning. It features complex backgrounds, multiple objects per image, and annotations for bounding boxes and segmentation masks. For our simulation, we used the COCO val2017 subset ( 5,000 images) to efficiently evaluate our quantized Faster R-CNN model on a diverse, well-annotated set, avoiding the computational cost of the full training set ( 118,000 images). Here, images are resized to $256\times 256$ followed by a crop to $224\times 224$ for analysis.
@@ 4.3.1. Image classifications
Figure 3. Image classification on Caltech101 using Custom(5, 5) precision.
Figure 4. Image classification on Caltech101 using bfloat16 precision.
Figure 5. Image classification on Caltech101 using tf32 precision.
Figure 6. Image classification on Caltech101 using fp32 precision.
@@ 4.3.2. Object detection
Table 9. mAP@0.5:0.95 by floating-point Representation and Rounding Method
@@
Figure 7. Object detection using bfloat16 precision (Red: Ground Truth, Green: Predictions).
Figure 8. Object detection tf32 precision (Red: Ground Truth, Green: Predictions).
Figure 9. Object detection using fp32 precision (Red: Ground Truth, Green: Predictions).
@@ 5. Conclusion and future work
In this work, we present the open-source software called Pychop, for efficient low-precision emulation for numerical methods and deep learning research. By integrating seamlessly with automatic differentiation frameworks such as PyTorch, JAX, and NumPy, Pychop enhance accessibility and usability in computational science developed by different scientific software. Its flexibility, customized design, and comprehensive rounding support establish it as a vital resource for advancing mixed-precision numerical algorithms and deep learning applications.
Empirical results across various rounding modes and precisions in Python and Matlab for low-precision floating-point emulation demonstrated that Pychop achieves a competing speedup over MATLAB chop, with improvements of multiple orders of magnitude when deployed on a GPU. Besides, we simulated post-quantization effects in image classification and object detection tasks on Caltech101, OxfordIIITPet, and COCO datasets to further validate its usage. These experiments offer valuable insights into the optimal bitwidths for exponents and significands needed to achieve high-quality inference, highlighting performance trade-offs to guide the selection of efficient floating-point representations tailored to specific use cases.
Pychop is engineered for sustained innovation and development. We plan to expand its compatibility with frameworks such as TensorFlow (Abadi et al., 2015) to increase its impact and adoption. We invite contributions from the research and development community to enhance functionality, optimize performance, and explore novel applications, ensuring Pychop remains competitive for future advancements.